OpenStudy (anonymous):

–7 + (–3)x + 5 + (–6)x

OpenStudy (anonymous):

since it's addition you can just simply remove the bracket and simplify..

OpenStudy (anonymous):

SIMPLIFY -7 + 5 =-2 -3x + -6x = -9x -2-9x